i.                Following the given equations of costs and its supposed price, find the equilibrium costs and equilibrium quantities

a.    ATC=3Q2 +12Q+350,                    P = 60

b.    TC = 300Q2 +10Q+120,          P = 40 



1
Expert's answer
2021-04-28T20:10:42-0400

(a)a. Given the ATC, calculate TC and the profit equation as below;

"ATC=3Q^{2 }+12Q+350"  

"ATC=\frac{TC}{Q}"

"3Q^{2} +12Q+350=\frac{TC}{Q}"

"TC=3Q^{3 }+12Q^{2}+350Q"

"Profits: \u03c0=PQ\u2212TC"

"\u03c0=60Q\u22123Q^{3} \u221212Q^{2}\u2212350Q"

The firm will maximize profits w.r.t. Q;

At profit maximizing level of Q;

"\frac{∂π}{∂Q}=0"

"60\u22129Q^{2} \u221224Q\u2212350=0"

"9Q^{2} +24Q+290=0"

The above quadratic equation in Q has a negave determinant. 

The equation is not satified for any positive 

value of Q.

Q=0


(b)"TC=300Q^{2}+10Q+120"

"\frac{dTC}{dQ}=MC=600Q+10"

"TR=P\u00d7Q"

"=40Q"

"\frac{dTC}{dQ}=MR=40"

At equilibrium MR=MC

"40=600Q+10"

"Q=\frac{30}{600}"

"=0.05units"

At equilibrium cost

"TC=300(0.05)^{2}+10(0.05)+120"

"=0.75+0.5+120"

$"121.25"